OpenStudy (kewlgeek555):

I Need Someone to Just Explain to me How to Solve a Reclusive Formula and an Explicit Formula, maybe even give me pure English Definitions?

OpenStudy (kewlgeek555):

For example, I know that a reclusive formula would be: F(1) = 35, f(n) = f(1) + f(n – 1) for n > 1 but how do you solve it? I know that explicit formulas are: F(n) = 35n I think I know how to solve that...

OpenStudy (kewlgeek555):

@mathstudent55 @campbell_st @Firejay5 @agen0smith @DSS @savannaxx_ @Luigi0210 @DebbieG

OpenStudy (kewlgeek555):

@agent0smith

OpenStudy (agent0smith):

Not sure what you mean by "solve it"... the recursive formula gives you a way to find a term, given the previous term.

OpenStudy (kewlgeek555):

Oh...well. Hmm...let me see the instructions. Well, this one asks: b) Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations

OpenStudy (kewlgeek555):

@agent0smith the reclusive formula is F(1) = 35, f(n) = f(1) + f(n – 1) for n > 1

OpenStudy (debbieg):

The formula is recursive, not reclusive (although generally, it's true that formulas don't get out much.... lol ;) So you need the values of f(n) for n=1, 2, 3, 4, 5. You have f(1)=35 Now use the formula to get f(2): f(n) = f(1) + f(n – 1) f(2)= f(1) + f(2 – 1)= f(1) + f(1)=35 + 35 = 70 Now do the same thing for n=3, 4, and 5. There are your values.